# Clever: A Curated Benchmark for Formally Verified Code Generation

### v3

@@ 1 Introduction
To date, there are a handful of benchmarks [9, 27, 26] for formally verified code generation. However, the formal specifications in these benchmarks tend not to capture the full (natural-language) intent behind the target program and sometimes hint at ways to implement the program. This ambiguity allows a code generator to “cheat” by generating trivial programs or copying code from the specification (see Section A.1).
@@ 2 The Clever Benchmark
These steps (Figure 1) form two certification goals: (1) Specification certification: Steps 1–2 verify that the model correctly inferred the intended behavior. (2) Implementation certification: Steps 3–4 verify that the generated implementation satisfies the formal intent.
@@
{mdframed}
[roundcorner=10pt] (a) -- computable spec def problem_spec -- function signature (implementation: Nat Nat) -- inputs (n: Nat) : Prop := -- spec let spec (result: Nat) := match n with | 0 => result = 0 | 1 => result = 1 | n + 2 => result = implementation n + implementation (n + 1) -- return value satisfies spec result, implementation n = result spec result (b) -- non-computable spec inductive fibonacci_non_computable : Prop | base0 : fibonacci_non_computable 0 0 | base1 : fibonacci_non_computable 1 1 | step : n f f₂, fibonacci_non_computable n f fibonacci_non_computable (n + 1) f fibonacci_non_computable (n + 2) (f + f₂) def problem_spec -- function signature (implementation: Nat Nat) -- inputs (n: Nat) := -- spec let spec (result: Nat) := fibonacci_non_computable n result -- program termination result, implementation xs = result spec result
@@
A key design decision in our benchmark is the use of non-computable specifications, which are predicates or functions in Lean that return propositions (Prop in Lean) that cannot be evaluated or simplified (decided by Lean) through computation alone. These contrast with computable specifications, written as executable functions or decidable predicates that Lean can reduce directly. While easier to verify, computable specs often leak the desired logic: models can copy them into implementations and produce trivial proofs via rewriting. Figure 2 shows the difference between a computable and a non-computable specification.
Figure 3 demonstrates the importance of this contrast. The left side (a–c) shows a computable spec whose logic is mirrored exactly in the GPT-4o-generated implementation, enabling a trivial proof. On the right (d–f), the spec is non-computable and requires symbolic reasoning to prove correctness. Notably, the GPT-4o-generated implementation in (e) does not mirror the spec, and the proof fails without further reasoning. This design ensures that models must engage in deeper logical inference, not just syntactic pattern matching. By using non-computable specs across our benchmark, we eliminate leakage and enforce truly verified reasoning from models.
{mdframed}
[roundcorner=6pt, innerleftmargin=4pt, innerrightmargin=4pt, innertopmargin=2pt, innerbottommargin=2pt] (a) def problem_spec (implementation: List Int Int Bool) (q: List Int) (w: Int) := let spec (result : Bool) := result (List.Palindrome q) (List.sum q w) result, implementation q w = result spec result (b) def implementation (q: List Int) (w: Int) : Bool := -- implementation generated by GPT-4o List.Palindrome q List.sum q w (c) theorem correctness (q: List Int) (w: Int) : problem_spec implementation q w := by -- proof generated by GPT-4o unfold problem_spec let result := implementation q w use result simp [result] simp [implementation] (d) def problem_spec (implementation: List Int Int Bool) (q: List Int) (w: Int) := let spec (result : Bool) := (result (List.Palindrome q)) (result (List.sum q w)) (¬(List.Palindrome q) ¬ result) (¬(List.sum q w) ¬ result) result, implementation q w = result spec result (e) def implementation (q: List Int) (w: Int) : Bool := -- implementation generated by GPT-4o let is_palindrome := q = q.reverse let sum_le_w := q.sum w is_palindrome && sum_le_w (f) theorem correctness (q: List Int) (w: Int) : problem_spec implementation q w := by -- proof generated by GPT-4o unfold problem_spec let result := implementation q w use result simp [result] simp [implementation] intro h -- <- The compilation fails here simp [h] exact List.eq_reverse_of_palindrome h.left -- more proof trimmed
@@
Curating the benchmark also revealed deeper challenges inherent to formal verification. For instance, in the HumanEval problem involving root-finding for polynomials (see Figure 4), proving termination is difficult due to reliance on unbounded numerical search. Similarly, generating verified code for “finding all prime Fibonacci numbers” encounters foundational roadblocks, as there is no known proof that infinitely many such numbers exist—highlighting how natural language tasks can conceal deep mathematical issues when formalized. One potential way to deal with these types of formulations is by adding the concept of computational fuel and approximate answers (see Figure 4, and Figure 8 in Section A.2). Writing non-computable specifications is particularly challenging for problems that rely on language-level features like Python’s eval, as seen in Problem 160. Since Lean lacks direct string-based evaluation, we had to reconstruct the behavior using inductive definitions over token lists and arithmetic expressions. This required converting a naturally computable task into a semantically equivalent, non-computable formulation without leaking implementation details. As shown in Figure 10 (in Section A.3), achieving this often involves layered recursive structures and careful abstraction to ensure both correctness and opacity.
{mdframed}
@@
[roundcorner=10pt, innerleftmargin=6pt, innerrightmargin=6pt, innertopmargin=4pt, innerbottommargin=4pt] (a) def problem_spec -- function signature (implementation: List Rat Rat) -- inputs (xs: List Rat) := -- spec let spec (result: Rat) := let eps := (1: Rat) / 1000000; xs.length 1 xs.length % 2 = 0 poly : Polynomial Rat, poly.degree = some (xs.length - 1) (∀ i, i xs.length - 1 poly.coeff i = xs.get! i) |poly.eval result| eps; -- program termination result, implementation xs = result spec result (b) -- possible implementation using Newtons method def implementation (xs: List Rat) : Rat := let rec poly (xs: List Rat) (x: Rat) := xs.reverse.foldl acc a => acc * x + a) 0; let rec poly (xs: List Rat) (x: Rat) := (xs.drop 1).reverse.foldl acc a => acc * x + a) 0; let rec eps := (1: Rat) / 1000000; let rec find_zero (xs: List Rat) (guess: Rat) (fuel: Nat) := let eval := poly xs guess; let eval := poly xs guess; if eval eps fuel = 0 then (guess, fuel) else let guess := (eval * guess - eval) / eval’; find_zero xs guess (fuel - 1); (find_zero xs 1.0 1000000).1
@@
Figure 4: Polynomial Root-Finding. Problem 32 asks for an approximate real root of a degree-$n$ polynomial. The spec enforces small residual error ($<10^{-6}$). The implementation uses Newton’s method with bounded recursion; proving termination is non-trivial due to lack of guaranteed derivative behavior.
@@
Another instructive case is the problem of computing the MD5 checksum (problem 162). Here, the formal specification must, by necessity, describe the exact computation, making it closely related to the implementation itself. Since we could not find any popular hashing libraries in Lean, we chose not to formalize this specific problem. However, we prescribe the recipe for creating non-computable definitions in Section A.3, given that we know the computable definition.
@@
Prior work, such as FVAPPS [9], relies on automatically generated specifications that can be incomplete or leaky, allowing trivial implementations (e.g., always returning zero) to pass (see Figure 7 in Section A.1). Our human-curated specifications ensure completeness and robustness, closing such loopholes and surfacing the real verification complexity hidden in everyday programming problems.
@@ 3 Evaluation
To fairly compare approaches that differ in model size, latency, and API usage, we adopt the metric pass@k-seconds—the fraction of benchmark problems solved within a fixed time budget $k$. A task is marked as solved only if both the formal specification and the implementation are generated and verified via Lean’s type checker. As described in Figure 5, each step in the Clever pipeline (spec generation, equivalence proof, implementation, and correctness proof) is retried until a valid Lean-compilable output is found or the time runs out.
[…]
@@
Results.
In particular, implementation compilation includes not only type-checking against the declared function signature, but also validation against a suite of example-based test cases adapted from the original HumanEval prompts. While passing these tests provides some evidence of functional correctness[25], we deliberately exclude them from our core success metric—since test cases offer only partial coverage and cannot guarantee semantic soundness (see Section 2 for discussion).
As shown in Table 1, compilation rates are broadly similar across few-shot models for both specification and implementation generation. A notable exception is the higher implementation compilation rate achieved by o4-mini, which contrasts with its lower success in proving correctness. More generally, even when an approach successfully certifies multiple specifications or verifies correctness for multiple implementations, the overall end-to-end success rate remains low. This is largely due to mismatch: tasks for which specification certification is tractable are often those where implementation correctness proofs are especially difficult, and vice versa. As a result, the joint success condition is rarely satisfied.
@@ Results.
Table 1: Evaluation of different strategies for end-to-end verified code generation. Each approach consists of five components: Model (LLM used), Spec Gen (formal specification generation), Equiv Proof (proof of equivalence to ground-truth spec), Impl Gen (program synthesis), and Corr Proof (proof of implementation correctness). FS indicates few-shot prompting with 1–2 examples. Evaluation follows the pipeline in Figure 5. Pass@k-seconds with $k=600$ reports the fraction of tasks where Lean successfully compiles the outputs and accepts the associated proofs within a 600-second time budget. The Compiled columns indicate whether the generated Lean code is syntactically valid and type-checks. The Proved columns reflect whether the corresponding proofs were accepted by Lean’s kernel, thereby certifying semantic correctness. The End-to-End column reports full pipeline success—i.e., both the specification and implementation must compile and their respective proofs must be accepted. Despite strong models like GPT-4o achieving high compilation rates, formal correctness remains challenging: no approach has yet succeeded across all stages on more than one problem (specifically problem 53).
@@
As shown in Table 2, proofs for specification certification are consistently longer and harder to generate than those for implementation correctness. This is expected: proving that a generated spec is semantically equivalent to a non-computable reference specification requires models (or agents) to reason abstractly about intent, without access to implementation-level cues. In contrast, correctness proofs often benefit from direct pattern matching or automation through tactics like simp.
This distinction is especially evident in the only problem for which an end-to-end verified code generation succeeds across multiple models: problem 53, which asks for the sum of two integers. Despite the simplicity of the implementation, the ground-truth specification is expressed in a way that deliberately obfuscates the target behavior. This design makes the equivalence proof non-trivial and requires models (or COPRA) to recover the algebraic structure underlying addition. Even here, success is only possible because the proofs admit aggressive automation via simp and ring. The full problem is shown in Figure 6, which illustrates the separation between syntactic and semantic difficulty across spec, implementation, and proofs.
Notably, Claude-3.7 in combination with COPRA successfully solves every implementation certification task that any other approach is able to solve. Figure 21 in Section A.5 illustrates one such case, showcasing a 35-line proof for the Brazilian factorial task that requires symbolic reasoning over factorial identities and recursive structure.
@@
Table 2: Analysis of successfully generated proofs across different models and certification types. We report: (1) the number of problems for which the correctness (isomorphism resp.) proofs are found by the approach in the column “# Qed” (see Figure 1), (2) the average number of lines in the proof, (3) the range of proof lengths (min–max), and (4) the average time it took for the approach to find a proof (given a proof was found). This analysis highlights variation in proof complexity and model behavior across settings. Few-shot prompting typically yields shorter, more brittle proofs, while COPRA-augmented configurations show higher robustness, with more consistent success and a broader range of proof strategies. Proof line counts serve as a coarse indicator of reasoning complexity.
{mdframed}
[roundcorner=6pt, innerleftmargin=6pt, innerrightmargin=6pt, innertopmargin=4pt, innerbottommargin=4pt]
def problem_spec (impl : Int Int Int) (x y : Int) :=
let spec (res : Int) := res - x - y = 0
result, impl x y = result spec result
def generated_spec (impl : Int Int Int) (x y : Int) : Prop :=
impl x y = x + y
def implementation (x y : Int) : Int := x + y
theorem correctness (x y : Int) : problem_spec implementation x y :=
by
unfold problem_spec
let result := implementation x y
use result
simp [result]
simp [implementation]
theorem spec_isomorphism :
impl, (∀ x y, problem_spec impl x y)
(∀ x y, generated_spec impl x y) :=
by
intro impl
apply Iff.intro
-- direction
intro h_prob_spec
intro x y
have h := h_prob_spec x y
simp [generated_spec, problem_spec] at h
rw [generated_spec]
linarith
-- direction
intro h_gen_spec
intro x y
unfold problem_spec
simp
have h := h_gen_spec x y
simp [generated_spec] at h
rw [h]
ring

### v4

@@ 1 Introduction
To date, there are a handful of benchmarks [9, 27, 26] for formally verified code generation. However, the formal specifications in these benchmarks tend not to capture the full (natural-language) intent behind the target program and sometimes hint at ways to implement the program. This ambiguity allows a code generator to “cheat” by generating trivial programs or copying code from the specification (see Section˜A.1).
@@ 2 The Clever Benchmark
These steps (Figure˜1) form two certification goals: (1) Specification certification: Steps 1–2 verify that the model correctly inferred the intended behavior. (2) Implementation certification: Steps 3–4 verify that the generated implementation satisfies the formal intent.
@@
(a) -- computable spec def problem_spec -- function signature (implementation: Nat Nat) -- inputs (n: Nat) : Prop := -- spec let spec (result: Nat) := match n with | 0 => result = 0 | 1 => result = 1 | n + 2 => result = implementation n + implementation (n + 1) -- return value satisfies spec result, implementation n = result spec result (b) -- non-computable spec inductive fibonacci_non_computable : Prop | base0 : fibonacci_non_computable 0 0 | base1 : fibonacci_non_computable 1 1 | step : n f f₂, fibonacci_non_computable n f fibonacci_non_computable (n + 1) f fibonacci_non_computable (n + 2) (f + f₂) def problem_spec -- function signature (implementation: Nat Nat) -- inputs (n: Nat) := -- spec let spec (result: Nat) := fibonacci_non_computable n result -- program termination result, implementation xs = result spec result
@@
A key design decision in our benchmark is the use of non-computable specifications, which are predicates or functions in Lean that return propositions (Prop in Lean) that cannot be evaluated or simplified (decided by Lean) through computation alone. These contrast with computable specifications, written as executable functions or decidable predicates that Lean can reduce directly. While easier to verify, computable specs often leak the desired logic: models can copy them into implementations and produce trivial proofs via rewriting. Figure˜2 shows the difference between a computable and a non-computable specification.
Figure˜3 demonstrates the importance of this contrast. The left side (a–c) shows a computable spec whose logic is mirrored exactly in the GPT-4o-generated implementation, enabling a trivial proof. On the right (d–f), the spec is non-computable and requires symbolic reasoning to prove correctness. Notably, the GPT-4o-generated implementation in (e) does not mirror the spec, and the proof fails without further reasoning. This design ensures that models must engage in deeper logical inference, not just syntactic pattern matching. By using non-computable specs across our benchmark, we eliminate leakage and enforce truly verified reasoning from models.
(a) def problem_spec (implementation: List Int Int Bool) (q: List Int) (w: Int) := let spec (result : Bool) := result (List.Palindrome q) (List.sum q w) result, implementation q w = result spec result (b) def implementation (q: List Int) (w: Int) : Bool := -- implementation generated by GPT-4o List.Palindrome q List.sum q w (c) theorem correctness (q: List Int) (w: Int) : problem_spec implementation q w := by -- proof generated by GPT-4o unfold problem_spec let result := implementation q w use result simp [result] simp [implementation] (d) def problem_spec (implementation: List Int Int Bool) (q: List Int) (w: Int) := let spec (result : Bool) := (result (List.Palindrome q)) (result (List.sum q w)) (¬(List.Palindrome q) ¬ result) (¬(List.sum q w) ¬ result) result, implementation q w = result spec result (e) def implementation (q: List Int) (w: Int) : Bool := -- implementation generated by GPT-4o let is_palindrome := q = q.reverse let sum_le_w := q.sum w is_palindrome && sum_le_w (f) theorem correctness (q: List Int) (w: Int) : problem_spec implementation q w := by -- proof generated by GPT-4o unfold problem_spec let result := implementation q w use result simp [result] simp [implementation] intro h -- <- The compilation fails here simp [h] exact List.eq_reverse_of_palindrome h.left -- more proof trimmed
@@
Curating the benchmark also revealed deeper challenges inherent to formal verification. For instance, in the HumanEval problem involving root-finding for polynomials (see Figure˜4), proving termination is difficult due to reliance on unbounded numerical search. Similarly, generating verified code for “finding all prime Fibonacci numbers” encounters foundational roadblocks, as there is no known proof that infinitely many such numbers exist—highlighting how natural language tasks can conceal deep mathematical issues when formalized. One potential way to deal with these types of formulations is by adding the concept of computational fuel and approximate answers (see Figure˜4, and Figure˜9 in Section˜A.2). Writing non-computable specifications is particularly challenging for problems that rely on language-level features like Python’s eval, as seen in Problem 160. Since Lean lacks direct string-based evaluation, we had to reconstruct the behavior using inductive definitions over token lists and arithmetic expressions. This required converting a naturally computable task into a semantically equivalent, non-computable formulation without leaking implementation details. As shown in Figure˜11 (in Section˜A.3), achieving this often involves layered recursive structures and careful abstraction to ensure both correctness and opacity.
@@
(a) def problem_spec -- function signature (implementation: List Rat Rat) -- inputs (xs: List Rat) := -- spec let spec (result: Rat) := let eps := (1: Rat) / 1000000; xs.length 1 xs.length % 2 = 0 poly : Polynomial Rat, poly.degree = some (xs.length - 1) (∀ i, i xs.length - 1 poly.coeff i = xs.get! i) |poly.eval result| eps; -- program termination result, implementation xs = result spec result (b) -- possible implementation using Newtons method def implementation (xs: List Rat) : Rat := let rec poly (xs: List Rat) (x: Rat) := xs.reverse.foldl acc a => acc * x + a) 0; let rec poly (xs: List Rat) (x: Rat) := (xs.drop 1).reverse.foldl acc a => acc * x + a) 0; let rec eps := (1: Rat) / 1000000; let rec find_zero (xs: List Rat) (guess: Rat) (fuel: Nat) := let eval := poly xs guess; let eval := poly xs guess; if eval eps fuel = 0 then (guess, fuel) else let guess := (eval * guess - eval) / eval’; find_zero xs guess (fuel - 1); (find_zero xs 1.0 1000000).1
@@
Figure 4: Polynomial Root-Finding. Problem 32 asks for an approximate real root of a degree-$n$ polynomial. The spec enforces small residual error ($10^{-6}$). The implementation uses Newton’s method with bounded recursion; proving termination is non-trivial due to lack of guaranteed derivative behavior.
@@
Another instructive case is the problem of computing the MD5 checksum (problem 162). Here, the formal specification must, by necessity, describe the exact computation, making it closely related to the implementation itself. Since we could not find any popular hashing libraries in Lean, we chose not to formalize this specific problem. However, we prescribe the recipe for creating non-computable definitions in Section˜A.3, given that we know the computable definition.
@@
Prior work, such as FVAPPS [9], relies on automatically generated specifications that can be incomplete or leaky, allowing trivial implementations (e.g., always returning zero) to pass (see Figure˜8 in Section˜A.1). Our human-curated specifications ensure completeness and robustness, closing such loopholes and surfacing the real verification complexity hidden in everyday programming problems.
@@ 3 Evaluation
To fairly compare approaches that differ in model size, latency, and API usage, we adopt the metric pass@k-seconds—the fraction of benchmark problems solved within a fixed time budget $k$. A task is marked as solved only if both the formal specification and the implementation are generated and verified via Lean’s type checker. As described in Figure˜5, each step in the Clever pipeline (spec generation, equivalence proof, implementation, and correctness proof) is retried until a valid Lean-compilable output is found or the time runs out.
We also compare the different approaches for each task using pass@k[4] (compile@k and prove@k respectively).
  • $\textnormal{Evaluate}(approach,\ \text{timeout})$
@@
Hybrid Baseline
GPT-5-mini FS Kimina FS FS Kimina FS 90.062% 0% 84.472% 0.621% 0%
Table 1: Evaluation of different strategies for end-to-end verified code generation. Each approach consists of five components: Model (LLM used), Spec Gen (formal specification generation), Equiv Proof (proof of equivalence to ground-truth spec), Impl Gen (program synthesis), and Corr Proof (proof of implementation correctness). FS indicates few-shot prompting with 1–2 examples. Evaluation follows the pipeline in Figure˜5. Pass@k-seconds with $k=600$ reports the fraction of tasks where Lean successfully compiles the outputs and accepts the associated proofs within a 600-second time budget. The Compiled columns indicate whether the generated Lean code is syntactically valid and type-checks. The Proved columns reflect whether the corresponding proofs were accepted by Lean’s kernel, thereby certifying semantic correctness. The End-to-End column reports full pipeline success—i.e., both the specification and implementation must compile and their respective proofs must be accepted. Despite strong models like GPT-4o achieving high compilation rates, formal correctness remains challenging: no approach has yet succeeded across all stages on more than one problem (specifically problem 53).
Results.
(a) Specification Compilation (compile@k)
(b) Specification Certification (prove@k)
(c) Implementation Compilation (compile@k)
(d) Implementation Certification (prove@k)
Figure 6: Aggregated compile@k and prove@k results across $k$ attempts, diagnosing failure modes for specification and implementation generation. Dotted lines indicate extrapolation of the pass rate beyond the 600s timeout. (a) Specification Compilation: Most models achieve high compile@k rates, with Claude-3.7 and o4-mini reaching >80%. (b) Specification Certification: Proving specification equivalence is a major bottleneck. All few-shot models solve only one problem (0.62%), while GPT-4o (CoPRA) proves  1.8%. (c) Implementation Compilation: o4-mini has the highest compilation rate at >80%, while other models cluster between 60-70%. (d) Implementation Certification: Claude-3.7 (CoPRA) performs best, certifying 8.7% of implementations, followed by DeepSeek-R1 (Few-shot) at 5.6%. These plots show that while models are fluent at generating compilable artifacts, formal certification remains the key challenge, with all prove@k rates below 10%.
In particular, implementation compilation includes not only type-checking against the declared function signature, but also validation against a suite of example-based test cases adapted from the original HumanEval prompts. While passing these tests provides some evidence of functional correctness[25], we deliberately exclude them from our core success metric—since test cases offer only partial coverage and cannot guarantee semantic soundness (see Section˜2 for discussion).
As shown in Figure˜6 and Table˜1, compilation rates are broadly similar across few-shot models for both specification (Figure˜6(a)) and implementation (Figure˜6(c)) generation. A notable exception is the higher implementation compilation rate achieved by o4-mini, which contrasts with its lower success in proving correctness. More generally, even when an approach successfully certifies multiple specifications (Figure˜6(b)) or verifies correctness for multiple implementations (Figure˜6(d)), the overall end-to-end success rate remains low. This is largely due to mismatch: tasks for which specification certification is tractable are often those where implementation correctness proofs are especially difficult, and vice versa. As a result, the joint success condition is rarely satisfied.
Finally, we evaluated KiminaProver-7b[33], a specialized prover. Due to its non-standard output formatting which posed parsing challenges, we used it only for proof steps, pairing it with GPT-5-mini for generation. This hybrid (Table˜1) yielded the highest compilation rates (90.1% spec, 84.5% impl), highlighting GPT-5-mini’s fluency. However, it certified almost nothing (0 spec, 1 impl), as the pipeline was bottlenecked by these same parsing issues.
@@ Results.
As shown in Table˜2, proofs for specification certification are consistently longer and harder to generate than those for implementation correctness. This is expected: proving that a generated spec is semantically equivalent to a non-computable reference specification requires models (or agents) to reason abstractly about intent, without access to implementation-level cues. In contrast, correctness proofs often benefit from direct pattern matching or automation through tactics like simp.
This distinction is especially evident in the only problem for which an end-to-end verified code generation succeeds across multiple models: problem 53, which asks for the sum of two integers. Despite the simplicity of the implementation, the ground-truth specification is expressed in a way that deliberately obfuscates the target behavior. This design makes the equivalence proof non-trivial and requires models (or COPRA) to recover the algebraic structure underlying addition. Even here, success is only possible because the proofs admit aggressive automation via simp and ring. The full problem is shown in Figure˜7, which illustrates the separation between syntactic and semantic difficulty across spec, implementation, and proofs.
Notably, Claude-3.7 in combination with COPRA successfully solves every implementation certification task that any other approach is able to solve. Figure˜22 in Section˜A.5 illustrates one such case, showcasing a 35-line proof for the Brazilian factorial task that requires symbolic reasoning over factorial identities and recursive structure.
@@
Table 2: Analysis of successfully generated proofs across different models and certification types. We report: (1) the number of problems for which the correctness (isomorphism resp.) proofs are found by the approach in the column “# Qed” (see Figure˜1), (2) the average number of lines in the proof, (3) the range of proof lengths (min–max), and (4) the average time it took for the approach to find a proof (given a proof was found). This analysis highlights variation in proof complexity and model behavior across settings. Few-shot prompting typically yields shorter, more brittle proofs, while COPRA-augmented configurations show higher robustness, with more consistent success and a broader range of proof strategies. Proof line counts serve as a coarse indicator of reasoning complexity.
(a) def problem_spec (impl : Int Int Int) (x y : Int) := let spec (res : Int) := res - x - y = 0 result, impl x y = result spec result (b) def generated_spec (impl : Int Int Int) (x y : Int) : Prop := impl x y = x + y (c) def implementation (x y : Int) : Int := x + y (d) theorem correctness (x y : Int) : problem_spec implementation x y := by unfold problem_spec let result := implementation x y use result simp [result] simp [implementation] (e) theorem spec_isomorphism : impl, (∀ x y, problem_spec impl x y) (∀ x y, generated_spec impl x y) := by intro impl apply Iff.intro -- direction intro h_prob_spec intro x y have h := h_prob_spec x y simp [generated_spec, problem_spec] at h rw [generated_spec] linarith -- direction intro h_gen_spec intro x y unfold problem_spec simp have h := h_gen_spec x y simp [generated_spec] at h rw [h] ring